OpenStudy (anonymous):

Simplify the rational expressions. State any excluded values. Show your work. 3x-6/x-2

OpenStudy (anonymous):

3x-6=3(x-2) now you can Simplify this.

OpenStudy (amorfide):

factorise the top by taking out a factor of 3x and -6 the largest factor is 3 3(x-2) now you have \[\frac{ 3(x-2) }{ (x-2) }\] the (x-2)/(x-2) become 1 so 3(1)=3 x can not be equal to 2 as this would divide by 0 \[x \neq 2\]